# Finite Sample Analysis of Distribution-Free Confidence Ellipsoids for Linear Regression

### v1

@@ V Simulation Experiments
In this section we illustrate the sizes of the SPS-Indicator and SPS ellipsoidal outer approximation (EOA) confidence regions with simulation experiments, furthermore, we compare them with our theoretical bounds. Throughout the experiments, we considered a $2$-dimensional linear regression problem, cf. (1), with $\varphi_{t,i}\sim\text{Unif}(1,2)$ and $W_{t}\sim\text{Unif}(-3,3)$. Hence, the noises were symmetric, nonatomic and $\sigma$-subgaussian with $\sigma=3$. The confidence regions were generated with hyper-parameters $m=100$ and $q=50$, i.e., their significance level was $0.5$. We set the sample size to $n=2000$ and repeated the confidence region constructions for $s=100$ independently simulated trajectories. The parameters that are needed for the given theoretical bounds were computed as follows: $\kappa$ was the largest empirical value of $\frac{t}{d}\max_{1\leq i\leq t}\left\lVert\Phi_{Q,t}^{\mathrm{T}}e_{i}\right\rVert^{2}$, while $\lambda_{0}$ was the smallest empirical eigenvalue of $\bar{R}_{t}$ over all $t_{0}\leq t\leq n$ and all simulated trajectories. Notice that a consequence of setting the value of $\kappa$ this way is that $\rho=1$.
The empirical size of the SPS EOA region was computed as $2\sqrt{\gamma_{t}^{*}/\lambda_{\text{min}}(\bar{R}_{t})}$ for every $t_{0}\leq t\leq n$, since it corresponds to the longest axis of the ellipsoid $\sup_{\theta\in\mathcal{O}_{p,n}}2\hskip 0.85358pt\|\theta-\hat{\theta}_{n}\|$, see (12). In case of the SPS-Indicator regions, an extreme value of the region $\theta_{e}$ was obtained by doing a logarithmic search on the longest axis of the bounding ellipsoid, and then we used $2\hskip 0.85358pt\|\theta_{e}-\hat{\theta}_{n}\|$ as a numerical approximation of the diameter.
Figure 1: Comparison of the empirical sizes and the theoretical upper bounds on the sizes of the SPS-Indicator and SPS ellipsoidal outer approximation regions for $m=100$, $q=50$, $\delta=0.1$, $t_{0}=250$, $n=2000$ and $s=100$.
The difference between the empirical sizes of SPS-Indicator, SPS EOA and asymptotic confidence regions, furthermore, the theoretical bounds with confidence level $0.9$ ($\delta=0.1$) are shown in Fig. 1. We used quantiles for the empirical sizes in each iteration, i.e., the smallest number for which at least the specified portion $(s\cdot(1-\delta))$ of simulation realizations were below that number (from $s=100$ trials). The theoretical bounds were computed according to Theorems 2 and 3.
@@
The results of this numerical experiment illustrate that our theoretical bounds capture well the empirical decrease rates of the confidence regions as well as the difference between these rates on the sizes of SPS-Indicator and SPS EOA confidence regions. The results also indicate that our theoretical bounds are a bit conservative. This, however, is an expected phenomenon, since these empirical regions are built (a posteriori) with data-driven algorithms, while the theoretical bounds are calculated (a priori) from concentration inequalities.
VI Conclusion
@@ Appendix A Proof of Lemma 3
$\begin{cases}\dfrac{8d\sigma^{2}\ln^{\frac{1}{2}}(\tfrac{4}{\delta})}{n\lambda_{0}}\vspace{1mm}&4e^{-(nd\lambda_{0})^{2}}\leq\delta\leq 1,\\ \dfrac{8\sigma^{2}\ln(\tfrac{4}{\delta})}{n\lambda_{0}}&\delta<4e^{-(nd\lambda_{0})^{2}}.\end{cases}$ (70)
In the proof of Lemma 1 [20, Appendix B] it is shown that a random variable in the form of $X=w_{n}^{\mathrm{T}}Mw_{n}$, where $M$ is a projection matrix with rank($M$)$=d$, can be upper bounded as $X\leq\sum_{i=1}^{d}\tilde{w}_{d,i}^{2}$, where $\{\tilde{w}_{d,i}\}$ are (zero mean) $\sigma$-subgaussians. Since $X_{\scriptscriptstyle P}$ is in this form, we have that
@@
$\begin{cases}\dfrac{d\sigma^{2}\left(8\ln^{\frac{1}{2}}(\tfrac{4}{\delta})+1\right)}{n\lambda_{0}}\vspace{1mm}&4e^{-(nd\lambda_{0})^{2}}\leq\delta\leq 1,\\ \dfrac{\sigma^{2}\left(8\ln(\tfrac{4}{\delta})+d\right)}{n\lambda_{0}}&\delta<4e^{-(nd\lambda_{0})^{2}}.\end{cases}$ (73)
@@
$\begin{cases}\dfrac{d\sigma^{2}\left(8\ln^{\frac{1}{2}}(\tfrac{4}{\delta})+1\right)\left(n^{\frac{\rho}{2}}+g^{\frac{1}{2}}(\delta)\right)}{n\lambda_{0}\left(n^{\frac{\rho}{2}}-g^{\frac{1}{2}}(\delta)\right)}\vspace{1mm}&4e^{-(nd\lambda_{0})^{2}}\leq\delta\leq 1,\\ \dfrac{\sigma^{2}\left(8\ln(\tfrac{4}{\delta})+d\right)\left(n^{\frac{\rho}{2}}+g^{\frac{1}{2}}(\delta)\right)}{n\lambda_{0}\left(n^{\frac{\rho}{2}}-g^{\frac{1}{2}}(\delta)\right)}&\delta<4e^{-(nd\lambda_{0})^{2}}.\\ \end{cases}$
@@ Appendix B Technical Lemmas
In [25] it is showed that if the confidence region is bounded, then $A$ is positive definite, consequently $\bar{R}_{n}^{-\frac{1}{2}}A\bar{R}_{n}^{-\frac{1}{2}}$ is also positive definite (A2), therefore it is full rank. Using our definition of $A$ (25), it can be written that $\bar{R}_{n}^{-\frac{1}{2}}A\bar{R}_{n}^{-\frac{1}{2}}=\bar{R}_{n}^{-\frac{1}{2}}\Phi^{\mathrm{T}}D_{\alpha,n}(I-\Phi_{n}R_{n}^{-1}\Phi_{n}^{\mathrm{T}})D_{\alpha,n}\Phi\bar{R}_{n}^{-\frac{1}{2}}$. Lets denote $L\doteq\bar{R}_{n}^{-\frac{1}{2}}\Phi^{\mathrm{T}}D_{\alpha,n}(I-\Phi_{n}R_{n}^{-1}\Phi_{n}^{\mathrm{T}})$. Since $L$ is a part of the product of the $d$-ranked matrix $\bar{R}_{n}^{-\frac{1}{2}}A\bar{R}_{n}^{-\frac{1}{2}}$, it follows that rank$(L)=d$. Notice that $S_{1}(\hat{\theta}_{n})=1/n\cdot Lw_{n}$, therefore by applying Lemma 6 we conclude that $S_{1}(\hat{\theta}_{n})\neq 0$ almost surely, hence $\|S_{1}(\hat{\theta}_{n})\|^{2}\neq 0$ almost surely. ∎
@@
$BB^{\mathrm{T}}=\left(\Phi_{n}^{\mathrm{T}}-Q_{n}R_{n}^{-1}\Phi_{n}^{\mathrm{T}}D_{\alpha,n}\right)\left(\Phi_{n}-D_{\alpha,n}\Phi_{n}R_{n}^{-1}Q_{n}\right)=$
$\left(R_{n}-Q_{n}R_{n}^{-1}Q_{n}-Q_{n}R_{n}^{-1}Q_{n}+Q_{n}R_{n}^{-1}Q_{n}\right)=A.$ (88)

### v2

@@ V-A Sample complexity for bounded noise
The SPS EOA confidence regions were generated with hyper-parameters $m=10$ and $q=1$, i.e., their confidence level was $0.9$. The parameters that are needed for the sizes of SPS EOA theoretical bounds were computed as follows: $\kappa$ was the largest empirical value of $\frac{t}{d}\max_{1\leq i\leq t}\left\lVert\Phi_{Q,t}^{\mathrm{T}}e_{i}\right\rVert^{2}$, while $\lambda_{0}$ was the smallest empirical eigenvalue of $\bar{R}_{t}$ over all $t_{0}\leq t\leq n$ and all simulated trajectories. Notice that a consequence of setting the value of $\kappa$ this way is that $\rho=1$. The empirical size of the SPS EOA region was computed as $2\sqrt{\gamma_{t}^{*}/\lambda_{\text{min}}(\bar{R}_{t})}$ for every $t_{0}\leq t\leq n$, since it corresponds to the longest axis of the ellipsoid $\sup_{\theta\in\mathcal{O}_{p,n}}2\hskip 0.85358pt\|\theta-\hat{\theta}_{n}\|$, see (12). The sizes of the asymptotic and set membership confidence ellipsoids were computed accordingly.
We set the sample size to $n=2000$ and repeated the confidence region constructions for $s=100$ independently simulated trajectories. The difference between the empirical sizes of SPS EOA, set membership and asymptotic confidence regions, furthermore, the SPS EOA theoretical bounds with confidence level $0.5$ ($\delta=0.5$) are shown in Fig. 1. To reach the desired confidence level, the median was computed for the empirical sizes in each iteration from $s=100$ trials.
@@
The results of this numerical experiment illustrate that our theoretical bound capture well the empirical decrease rate of the SPS EOA confidence regions. The results also indicate that our theoretical bounds are a bit conservative. This, however, is an expected phenomenon, since these empirical regions are built (a posteriori) with data-driven algorithms, while the theoretical bounds are calculated (a priori) from concentration inequalities. The conservatism of the set membership region can be observed from this experiment, as well, even our theoretical bound has a smaller size for larger sample sizes. Also note that, while the asymptotic region has the smallest size, it lacks the strict theoretical guarantees that the other confidence sets provide, since it is only an approximate region.
Figure 2: Comparison of the $0.9$-level empirical and theoretical sizes of SPS EOA regions with asymptotic and DMR bound based ellipsoids for a non-stationary Gaussian mixture noise distribution with $m=10$, $q=1$, $\delta=0.5$, $t_{0}=400$, $n=2000$ and $s=100$.
V-B Sample complexity for unbounded noise
In a second experiment we compared our theoretical bound and the empirical sizes of SPS EOA and asymptotic regions with the finite sample DMR bound from [23, Theorem 3.1] given in Theorem 4. We considered a 2-dimensional FIR system with the same structure as in (71), but with a symmetric, non-stationary Gaussian mixture model given as
$W_{t}=\sum_{i=1}^{5}\mathbb{I}(\zeta=i)Z_{i,t},$ (73)
where $\zeta\in\{1,\dots,5\}$ has a discrete probability distribution: $\mathbb{P}(\zeta=1)=0.3$, $\mathbb{P}(\zeta=2)=\mathbb{P}(\zeta=3)=0.2$, $\mathbb{P}(\zeta=4)=\mathbb{P}(\zeta=5)=0.15$, and $Z_{1,t}\sim\mathcal{N}(0,(n-t)/n+1)$, $Z_{2,t}\sim\mathcal{N}(-2,3(n-t)/n)$, $Z_{3,t}\sim\mathcal{N}(2,3(n-t)/n)$, $Z_{4,t}\sim\mathcal{N}(-5,2(n-t)/n+1)$, $Z_{5,t}\sim\mathcal{N}(5,2(n-t)/n+1)$. This type of multimodal noise arises in various signal processing applications where distortions from different sources are modeled. Notice that the noise sequence is independent and symmetric, thus it satisfies the noise assumption on the SPS EOA construction, furthermore, it is $\sigma$-subgaussian with $\sigma=\sqrt{2.7}$. As the distribution is not identically distributed the asymptotic region can be only used as an approximation again. We also changed the input sequence w.r.t. the experiment in section V-A and removed the feedback from it to satisfy the assumption of Theorem 4, i.e., the input sequence was $U_{t}=\sum_{i=1}^{5}c_{i}V_{t-i+1}$ with the same parameters as before. From this setting it follows that $\sigma_{\varphi}=(\sum_{i=1}^{5}c_{i}^{2})^{1/2}$, and we choose the largest $C$ that satisfies (69) with $\nu=0.5$. Note that [23] does not specify the value of $C$ or how to chose the hyper-parameters of the bound, which makes it challenging to apply the DMR bound in practice.
We again considered $0.9$-level confidence regions, thus we set $m=10$, $q=1$ and $\eta=q/m$. The SPS EOA theoretical bound and the other confidence ellipsoids were computed the same way as before. Fig. 2 illustrates the difference between the empirical sizes of SPS EOA and asymptotic confidence regions with the theoretical DMR and SPS EOA size bounds.
It can be observed that initially (for small sample sizes) our bound on the SPS EOA sizes, given by Theorem 3, is more conservative than the DMR bound of Theorem 4. However, our bound decreases much faster as the sample size increases. This also shows that our bound is less conservative than another concentration inequality based uncertainty quantification method when the validity constraint ($n\geq\lceil g^{1/\rho}(\frac{\delta}{m-q})\rceil$) holds with a reasonable margin. Furthermore, it provides a theoretical guarantee for an algorithm that constructs much smaller confidence ellipsoids (here shown empirically) than the DMR bound. We emphasize again that although the asymptotic region is the smallest, its guarantees are only asymptotic, it can be over-optimistic.
V-C Higher-order FIR systems
TABLE I: Comparison of $0.5$-level confidence region sizes for higher-order FIR systems with $m=2$, $q=1$, $\delta=0.5$ and $s=100$.
d n SPS EOA SPS EOA bound DMR Asymptotic
4 1000 0.79 3.67 1.93 0.76
4 2000 0.55 1.47 1.92 0.53
6 2000 0.73 4.82 1.70 0.69
6 3500 0.53 1.61 1.69 0.51
8 3500 0.61 2.80 1.59 0.62
8 6000 0.45 1.33 1.58 0.44
10 6000 0.53 1.81 1.50 0.51
An experiment was also performed to investigate the sizes of SPS EOA, DMR-bound-based, and asymptotic confidence ellipsoids in higher-order FIR systems. Specifically, we studied
$Y_{t}=\sum_{k=1}^{d}b_{k}^{*}U_{t-k}+W_{t}.$ (74)
We considered $\forall\,k:b_{k}=5$, the same input sequence $U_{t}=\sum_{i=1}^{5}c_{i}V_{t-i+1}$ and a non-stationary Gaussian mixture noise sequence $\{W_{t}\}$, given in (73), as in the previous section. We generated the confidence ellipsoids with confidence level $0.5$, thus we set $m=2$ and $q=1$. The other parameters and the confidence ellipsoid sizes were computed as in Section V-B. The median of the empirical sizes from $s=100$ independent trials and the theoretical bounds for $\delta=0.5$ are presented in Table I for various sample sizes $n$ and dimensions $d$.
The results show similarity to the two dimensional case. Our theoretical bound provide more conservative sizes than the DMR bound when the sample size $n$ is close to the validity constraint ($n\geq\lceil g^{1/\rho}(\frac{\delta}{m-q})\rceil$), however, it can be seen that for a larger sample size, given the same dimension $d$, our bound is less conservative, i.e., it decreases faster. Table I also shows that the empirical sizes of SPS EOA regions is much smaller than of the DMR bounds and roughly the same as the asymptotic ones. As a result, the order of the FIR system (its dimension) does not influence significantly the relative sizes of the regions or the bounds when compared to each other.
V-D Computational Complexity
TABLE II: Comparison of the time required to solve the convex program relative to the $2$-dimensional case for $m=10$.
d n 250 500 750 1000
10 7.86 7.05 6.72 3.15
20 32.96 28.65 24.79 11.87
30 152.04 90.44 88.21 39.54
40 687.97 216.13 193.33 89.22
In our last experiment the computational complexity of calculating the SPS EOA confidence regions was studied. As one of the main advantages of SPS EOA regions is that they have rigorous guarantees even when there is just a limited number of measurements, we investigated the computational complexity for small sample sizes and relatively high dimensions. The main parameter that influences the computational time of the SPS EOA algorithm is the desired confidence level, since the convex program (27) has to be solved $m$ times. The time required to solve the convex programs in higher-dimensional problems relative to the $2$-dimensional case for different sample sizes and $m=10$ are shown in Table II.
It can be observed that when the dimension of the problem is relatively high compared to the number of observations, the time to solve the resulting SDP grows linearly. Nevertheless, the SPS EOA confidence regions can be computed efficiently on modern hardware, since the two-dimensional variant of the optimization problem can be solved in milliseconds. Furthermore, the results show that for relatively large $n/d$ ratios the convex problems can be solved considerably faster than for low ratios, showing the efficiency of the construction.
VI Discussion
Confidence regions, typically based on concentration inequalities, can provide high probability bounds on the estimation errors and have become one of the main tools for uncertainty quantification over the past decades. This field of study is important both in statistical learning and in engineering applications, where robustness is critical, such as adaptive control and filtering. Here, we focused on robust estimation of exogenous linear systems, including FIR models, which play a vital role in signal processing across various domains, e.g., communications, biomedical engineering, and radar systems.
One of the key advancement of this work is that we rigorously prove concentration inequality based high probability upper bounds on the sizes of SPS confidence ellipsoids, which are based on the solutions of convex semidefinite programs. The main advantages of SPS compared to the aforementioned methods is that it can build confidence regions without any specific distribution or moment assumptions in a data-driven fashion, i.e., without the need of any hyper-parameters. Our analysis provides rigorous guarantees for the sizes of these SPS ellipsoids, which can be compared to the result of [23]. Although it considers centered random regressor sequences, while our bound assumes deterministic regressors, the two bounds have the same decrease rate as the sample size increases and their dependence on the dimension and noise parameter is also similar. However, the DMR bound depends on hyper-parameters that are typically unknown in practice. Moreover, as we showed in our experiment for a robust FIR parameter estimation problem and a Gaussian input process, our bound gets less conservative than the DMR bound of [23] as the sample size increases. Furthermore, the empirical sizes of SPS EOA regions are much smaller than the DMR bound.
Therefore, as our finite sample PAC bound and the experimental evaluations demonstrate, SPS EOA has several advantages compared to concentration inequality based uncertainty quantification methods, such as a data-driven construction, smaller theoretical and significantly smaller empirical sizes.
One of the main assumptions behind SPS is that the noise terms are independent and symmetric about zero. A trade-off regarding this assumption can be made by using the Residual-Permuted Sums (RPS) algorithm [45], which builds confidence regions assuming that the noise terms are independent and identically distributed. A future research direction could be to analyze the sample complexity of RPS based confidence ellipsoids and compare them to the bounds of Theorems 3 and 4. An instrumental variable based extension of SPS to state-space models can be found in [34], while [46] proposed a generalization for ARX models. Another possible research direction could be to investigate the sizes of SPS confidence ellipsoids in case of closed-loop state space models or ARX systems and compare them with the results of [16, 22].
VII Conclusion
@@ Appendix A Proof of Lemma 3
$\begin{cases}\dfrac{8d\sigma^{2}\ln^{\frac{1}{2}}(\tfrac{4}{\delta})}{n\lambda_{0}}\vspace{1mm}&4e^{-(nd\lambda_{0})^{2}}\leq\delta\leq 1,\\ \dfrac{8\sigma^{2}\ln(\tfrac{4}{\delta})}{n\lambda_{0}}&0\leq\delta<4e^{-(nd\lambda_{0})^{2}}.\end{cases}$ (78)
In the proof of Lemma 1 [37, Appendix B] it is shown that a random variable in the form of $X=w_{n}^{\mathrm{T}}Mw_{n}$, where $M$ is a projection matrix with rank($M$)$=d$, can be upper bounded as $X\leq\sum_{i=1}^{d}\tilde{w}_{d,i}^{2}$, where $\{\tilde{w}_{d,i}\}$ are (zero mean) $\sigma$-subgaussians. Since $X_{\scriptscriptstyle P}$ is in this form, we have that
@@
$\begin{cases}\dfrac{d\sigma^{2}\left(8\ln^{\frac{1}{2}}(\tfrac{4}{\delta})+1\right)}{n\lambda_{0}}\vspace{1mm}&4e^{-(nd\lambda_{0})^{2}}\leq\delta\leq 1,\\ \dfrac{\sigma^{2}\left(8\ln(\tfrac{4}{\delta})+d\right)}{n\lambda_{0}}&0\leq\delta<4e^{-(nd\lambda_{0})^{2}}.\end{cases}$ (81)
@@
$\begin{cases}\dfrac{d\sigma^{2}\left(8\ln^{\frac{1}{2}}(\tfrac{4}{\delta})+1\right)\left(n^{\frac{\rho}{2}}+g^{\frac{1}{2}}(\delta)\right)}{n\lambda_{0}\left(n^{\frac{\rho}{2}}-g^{\frac{1}{2}}(\delta)\right)}\vspace{1mm}&4e^{-(nd\lambda_{0})^{2}}\leq\delta\leq 1,\\ \dfrac{\sigma^{2}\left(8\ln(\tfrac{4}{\delta})+d\right)\left(n^{\frac{\rho}{2}}+g^{\frac{1}{2}}(\delta)\right)}{n\lambda_{0}\left(n^{\frac{\rho}{2}}-g^{\frac{1}{2}}(\delta)\right)}&0\leq\delta<4e^{-(nd\lambda_{0})^{2}}.\\ \end{cases}$
@@ Appendix B Technical Lemmas
In [41] it is shown that if the confidence region is bounded, then $A$ is positive definite, consequently $\bar{R}_{n}^{-\frac{1}{2}}A\bar{R}_{n}^{-\frac{1}{2}}$ is also positive definite (A2), therefore it is full rank. Using our definition of $A$ (25), it can be written that $\bar{R}_{n}^{-\frac{1}{2}}A\bar{R}_{n}^{-\frac{1}{2}}=\bar{R}_{n}^{-\frac{1}{2}}\Phi^{\mathrm{T}}D_{\alpha,n}(I-\Phi_{n}R_{n}^{-1}\Phi_{n}^{\mathrm{T}})D_{\alpha,n}\Phi\bar{R}_{n}^{-\frac{1}{2}}$. Lets denote $L\doteq\bar{R}_{n}^{-\frac{1}{2}}\Phi^{\mathrm{T}}D_{\alpha,n}(I-\Phi_{n}R_{n}^{-1}\Phi_{n}^{\mathrm{T}})$. Since $L$ is a part of the product of the $d$-ranked matrix $\bar{R}_{n}^{-\frac{1}{2}}A\bar{R}_{n}^{-\frac{1}{2}}$, it follows that rank$(L)=d$. Notice that $S_{1}(\hat{\theta}_{n})=1/n\cdot Lw_{n}$, therefore by applying Lemma 6 we conclude that $S_{1}(\hat{\theta}_{n})\neq 0$ almost surely, hence $\|S_{1}(\hat{\theta}_{n})\|^{2}\neq 0$ almost surely. ∎
@@
$BB^{\mathrm{T}}$ $=\left(\Phi_{n}^{\mathrm{T}}-Q_{n}R_{n}^{-1}\Phi_{n}^{\mathrm{T}}D_{\alpha,n}\right)\left(\Phi_{n}-D_{\alpha,n}\Phi_{n}R_{n}^{-1}Q_{n}\right)$
$=\left(R_{n}-Q_{n}R_{n}^{-1}Q_{n}-Q_{n}R_{n}^{-1}Q_{n}+Q_{n}R_{n}^{-1}Q_{n}\right)$
$=A.$ (96)